OpenStudy (anonymous):

I give medals!!!!!!!!!!!!!!!!!! I just need to know how to do this! 8d-10 for d=4

OpenStudy (anonymous):

22

OpenStudy (anonymous):

8d-10 8(4)-10 32-10 the answer is 22

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

8(4)-10 = ? 32 - 10 = 22 just plug 4 in and you have the answer

OpenStudy (anonymous):

I forgot it says evaluate.

OpenStudy (anonymous):

Can you like type the instructions on how to do it. I don't necesarely want the answer, I just want to know how to do it.

OpenStudy (anonymous):

Insert d into where it says d. So (8 x 4) - 10 = 32 - 10 = 22

OpenStudy (anonymous):

Oh okay. Thanks!

OpenStudy (anonymous):

np